### 25 Dec 2023 v1

# Get a Grip: Reconstructing Hand-Object Stable Grasps in Egocentric Videos

Abstract
We address in-the-wild hand-object reconstruction for a known object category in egocentric videos, focusing on temporal periods of stable grasps. We propose the task of Hand-Object Stable Grasp Reconstruction (HO-SGR), the joint reconstruction of frames during which the hand is stably holding the object. We thus can constrain the object motion relative to the hand, effectively regularising the reconstruction and improving performance.
By analysing the 3D ARCTIC dataset, we identify temporal periods where the contact area between the hand and object vertices remain stable. We showcase that objects within stable grasps move within a single degree of freedom (1 DoF). We thus propose a method for jointly optimising all frames within a stable grasp by minimising the object’s rotation to that within a latent 1 DoF.
We then extend this knowledge to in-the-wild egocentric videos by labelling 2.4K clips of stable grasps from the EPIC-KITCHENS dataset. Our proposed EPIC-Grasps dataset includes 390 object instances of 9 categories, featuring stable grasps from videos of daily interactions in 141 environments. Our method achieves significantly better HO-SGR, both qualitatively and by computing the stable grasp area and 2D projection labels of mask overlaps.
1 Introduction
Accurately reconstructing three-dimensional hands along with a grasped or manipulated object is key to unlocking many perception problems, including fine-grained understanding of interactions, but also potential applications including augmented reality, imitation learning and human-machine interactions.
Prior approaches have laid the foundation for the task, though focusing mostly on 3D input [41, 2, 53], which limits its application to the vast amount of video footage. Approaches with 2D input have relied on 3D ground-truth for supervision [51, 7, 19, 23, 50], again restricting their scope to curated setups where multi-camera sensing and 3D ground truth can be obtained. Early insights [19, 5] demonstrate that evaluation on in-the-wild egocentric footage remains significantly challenging.
In this work, we simplify the task of reconstruction in-the-wild by focusing on temporal periods of stable grasps. Our contribution is composed of four components. First, we propose the task of Hand-Object Stable Grasp Reconstruction (HO-SGR) which jointly optimises the reconstructions across all frames within one stable grasp. Second, using the ARCTIC dataset [13], we showcase that objects move within one degree of freedom (DoF) once within the hand’s stable grasp. Third, we accordingly propose a method that jointly reconstructs the hands and objects by minimising the object’s motion to 1DoF across frames. We demonstrate our method outperforms baselines and alternative assumptions of object movement using 3D ground truth from the stable grasps within the egocentric views of ARCTIC. Fourth, we focus on our objective of reconstruction for in-the-wild egocentric videos. We label a sizable dataset of 2.4K stable grasps clips from egocentric videos. Our EPIC-Grasps dataset is the first for hand-object reconstruction collected from unscripted activities, with individuals grasping 390 different objects by both hands. Similar to previous works [17, 6, 15, 5], we restrict our evaluation to known category CAD models. Our dataset comes with pseudo-ground truth in the form of 2D segmentation masks available from [12], allowing to measure the 3D reconstruction’s projection relative to this 2D groundtruth.
We evaluate our proposed method on EPIC-Grasps, outperforming other baselines using the proxy measure of stable grasps within 2D projections. We demonstrate that joint optimisation within the stable grasp is vital in the egocentric setup as finger joints are often concealed from the camera viewpoint introducing additional ambiguity when reconstructing from a single image.
We demonstrate a sample from our EPIC-Grasps dataset along with reconstructions from our method in Fig LABEL:fig:teaser.
2 Related Works
For comparison of hand object reconstruction datasets, see Sec 5. We review related works in object pose estimation, hand-object reconstruction and grasping.
Object Pose Estimation The task of object pose estimation serves as a foundation for the broader task of hand-object reconstruction. Object pose estimation involves estimating the 6D poses of objects from given CAD models. Labbe and Michaud [27] and Wang et al. [44] proposed instance-level methods for object pose estimation. However, the NOCS method proposed by Wang et al. [45] assumes that no exact CAD models are available and requires an additional size estimation step. To address this limitation, Liu et al. [29] have proposed a method that estimates the 3-DoF sizes of objects, resulting in a 9-DoF pose estimation method.
In this work, we model the object reconstruction as estimating 6D poses with a scalar size, as in [19, 45].
3D Hand-Object-Reconstruction 3D hand object reconstruction methods can be grouped into two categories.
The first category, known-CAD methods [19, 5, 33], assumes that object CAD models are given and fits 3D shapes into 2D observations. To address the limited availability of 3D annotations, several methods [18, 28, 48] have been proposed. Other methods [46, 43, 50, 1] explore hand-object interactions through different modeling approaches. HOMan [19] is a generalisation of single-frame methods that incorporates temporal knowledge through the smoothness of mesh vertices over time and the propagation of object pose initialization using the previous frame’s pose.
The second category, CAD-agnostic methods, aims to estimate the hand and object poses without using explicit CAD models. An early attempt [42] models the object as a 3D bounding box and estimates its 6D pose. More recent methods, such as Diffusion-guided [52], ObMan [20] and implicit shape methods [51, 23, 7], use neural networks to encode the underlying object shape. However, recent methods like Huang et al. [22] require careful scanning of a firmly held object, making them impractical for use in in-the-wild videos.
Our dataset and method belong to the first category, which is a simplified assumption so we focus on the challenges of in-the-wild reconstruction.
3D Human Grasp Synthesis Generating plausible 3D human grasps for objects has been an active area of research.
Figure 1: We compare within/outside grasps, analysing object in-contact area (left), finger pose (middle) and relative object rotation w.r.t the hand (right), normalising all stable grasp duration for direct comparison (0 to 1 marked with red dotted lines). We showcase our manual annotations (top) and automatic (bottom).Objects do rotate within the stable grasp (blue) however mostly within 1 DoF (rotation residual in green).
GraspingField [23] characterises each point in a 3D space as the signed distances to the surfaces of the hand and the object, and learns a generative model to synthesize human grasps. ContactOpt [16] approximates deformable soft tissue in the hand by introducing a differentiable contact model. TOCH [56] learns to refine incorrect 3D hand-object interaction sequences using a correspondence-based prior. Zimeng et al. [55] utilises the physical stability to optimise single-image contact reconstruction.
However, as human grasp synthesis focuses on generating realistic grasps without incorporating input visual information, we do not explore these in this work.
3 What is a “stable grasp”?
The term “stable grasp” has been previously used in a number of prior works [4, 9, 14]. The term intuitively refers to a hand fully supporting the weight of the object, stopping it from slipping or sliding. During interactions, one hand holds an object in a “stable grasp” to either move it from one position to another, or for the second hand to manipulate the object. For example, if one is opening a bottle, then one hand is in a stable grasp while the other hand can be rotating the cap and is thus not in a stable grasp respect to the bottle.
Existing works on grasp type classification implicitly assume the given grasps are stable (or static) [14, 3, 35], without defining what a stable grasp is. Our first contribution is to explore what defines a stable grasp from a video sequence where the hand approaches an object, grasps it, firming the grip into a stable grasp and eventually releases the object. To do so, we leverage the recently introduced ARCTIC dataset [13]. The data offers a unique opportunity to study hand-object reconstruction as it includes 3D ground truth thanks to being captured in a MoCap system with 54 cameras and markers on both the hand and the objects.
For our exploration, we first analyse the footage and manually annotate 40 temporal extents of what we deem are stable grasps by visual inspection. Figure 1a shows that the stable grasp region corresponds to a consistent in-contact area during the temporal extent of the stable grasp—the same set of object vertices remain in contact with the hand during the grasp but change rapidly outside the grasp. We also demonstrate that both the finger tip poses and the object relative orientations do change during the stable grasps (average MPJPE [13] 4.6mm vs average rotation angle ${}^{\circ}$ for objects). However, these angles are significantly higher outside the stable grasp region (average MPJPE 11.8mm for fingers vs average rotation ${}^{\circ}$ for objects).
Figure 2: Sample stable grasp sequence from ARCTIC. Top: Dotted lines for start-end boundary. Bottom: stable contact area and object rotation along 1 DoF (purple).
Amongst the three metrics, we focus on the stable contact area as it gives the biggest change within/outside grasp. We thus define stable grasp as the temporal segment during which the object in-contact area has high intra-frame IOU
$\left[l^{*},r^{*}\right]=\underset{l,r}{\mathrm{argmax}}\left\{r-l\ :\mathrm{IOU}(S_{i},S_{j})>0.5\;\;\forall l\leq i<j\leq r\right\}$
where $S_{i}$ is the object vertices that are in-contact with the hand in frame $i$. That is the surface in-contact area should remain within an IOU of 0.5 to any other frames during the stable grasp. We use this definition to automatically find all stable grasp temporal boundaries in the ARCTIC dataset.
We accordingly locate 1303 stable grasp clips which we denote these as ARCTIC-Grasps dataset. In Fig 1b we compare this automatic set on the same metrics, showcasing consistent results. On average, 90% of frames have at least 74% IOU between any two frames in the stable grasp.
Object rotation within a stable grasp While object rotations within the stable grasp are consistently low (Fig 1 right), hands are so versatile that objects indeed do rotate, see Fig 2 for example. We observe that these rotations can be successfully approximated by rotations along a single degree-of-freedom (1 DoF). In Fig 1, we show the average rotation versus the rotation off a 1-DoF, calculated as the residual angle off this axis. This residual error is marginal within the stable grasp. In Sec 4.2 we elaborate on how we utilise this knowledge to optimise for hand object reconstructions within the stable grasp boundaries.
From 3D Datasets to Egocentric Videos. As noted in Sec 2 and our introduction, our objective is to perform HO-SGR in-the-wild without any 3D ground truth to assess the reconstruction. The absence of 3D ground truth has been an obstacle for evaluating performance in-the-wild, limiting it to qualitative results solely [52, 19]. By focusing on stable grasps, we offer a way forward to evaluation without 3D ground truth. We use the stable contact area as a measure of assessing successful reconstructions over the duration of the stable grasp. We then use the 1 DoF knowledge in HO-SGR, by minimising the rotation of the object off a latent axis of rotation, within which the object may still rotate. We explain our method next.
4 Method
Figure 3: Pipeline overview. We show 3 frames within a stable grasp. The yellow mesh color indicates 3D shapes that are being optimised; the blue mesh color indicates the underlying meshes are frozen. During the object relative pose optimisation stage, the object relative pose is optimised with 1-DoF constraint. During the hand-object refinement stage, the hand mesh and the object mesh are combined, and the ground-truth masks are merged.
4.1 Problem Statement and Notations
Given a sequence of images of hand-object interactions, annotated with hand-side, known object category, start-end times of stable grasp and 2D segmentation of the hand and object, we aim to reconstruct consistent grasping, in every sampled frame $n$ of the sequence of length $N$, by producing a per-frame pair of 3D meshes of the hand and the object w.r.t the camera. We refer to this task as Hand-Object Stable Grasp Reconstruction (HO-SGR).
To represent the hand mesh, we use the MANO [36] model, which takes as input a finger articulation vector $\theta\in R^{45}$ and an identity-specific vector $\beta\in\mathbb{R}^{10}$, and outputs the hand mesh with vertices $V_{h}\in\mathbb{R}^{778\times 3}$ in the hand coordinate system. We use FrankMocap [37] to obtain MANO finger articulations $\theta^{n}$ from all the frames and we assume the mean hand shape, i.e. $\beta^{n}=0$. We thus have per-frame hand vertices $V_{h}^{n}=\mathrm{MANO}(\theta^{n},\beta^{n})$. The hand-to-camera (h2c) pose $T_{h2c}^{n}\in SE(3)$ [32], which is defined as the hand wrist orientation and position, are produce by FrankMocap by default and are used to transform $V_{h}^{n}$ to $V_{h:c}^{n}$ in the camera coordinate system for each frame.
For the object mesh, we use $V_{o}\in\mathbb{R}^{|V_{o}|\times 3}$ to denote the known object vertices in the object coordinate system. We estimate the object-to-hand ($o2h$) poses $T_{o2h}^{n}\in SE(3)$ and the scale $s\in R$, which transform object vertices to $V_{o:h}^{n}$ in the hand coordinate system for each frame. We then use the hand-to-camera (h2c) pose $T_{h2c}^{n}$ to transform $V_{o:h}^{n}$ to $V_{o:c}^{n}$ in the camera coordinate system for each frame.
$V_{o:c}^{n}=T_{h2c}^{n}(T_{o2h}^{n}(s*V_{o}))$ (2)
4.2 Reconstructing Objects in a Stable Grasp
The stable grasp segments are defined based on the in-contact surface area, yet it is non-trivial to directly incorporate the Stable Contact Area constraint into optimisation, as the contact area is not known in advance.
Instead, as shown in Sec 3, we constrain the reconstruction so the in-hand object motion is minimised to a single degree-of-freedom, i.e. the object rotates along a fixed latent axis. We represent the 6D object-to-hand transform as a composition of transformation functions
$T_{o2h}^{n}=P_{o2h}\circ P^{-1}_{o2\phi}\circ R_{\phi}^{n}\circ P_{o2\phi}$ (3)
i.e. a per-frame rotation $R_{\phi}^{n}$ is first applied to the object along the latent axis, defined by $P_{o2\phi}$, then the common object-to-hand transform $P_{o2h}$ is applied. Furthermore, this does not restrict the ability of the hand to move relative to the camera or the wrist to move freely around the joint.
By back-propagating the error through $V_{o:h}^{n}$, the common parameters across all frames ($P_{o2h}$, $P_{o2\phi}$, $s$) as well as the individual frame parameters $\{R_{\phi}^{n}\}$ will be optimised jointly.
4.3 Initial Hand Mesh Refinement
Before optimising the object pose, we perform two steps of refinements for the FrankMocap hand output: temporal smoothness and 2D hand mask. We describe these next.
We define a general temporal smoothness function:
$F_{s}(\{\boldsymbol{x}^{n}\})=\sum_{n=2}^{N-1}|\boldsymbol{x}^{n}-\frac{\boldsymbol{x}^{n-1}+\boldsymbol{x}^{n+1}}{2}|^{2}_{2}$ (4)
We first optimise the finger articulation vector $\theta^{n}$ with the above temporal smoothness objective:
$E(\{\theta^{n}\})=F_{s}(\{\theta^{n}\})$ (5)
Then, we use the projected 2D hand mask to optimise the hand pose at every frame. We optimise the hand-to-camera pose $T_{h2c}^{n}=[R_{h}^{n}|t_{h}^{n}]$ for the hand mask pixel difference with ground-truth as well as temporal smoothness of poses:
$E(\{R_{h}^{n}\},\{t_{h}^{n}\})$ $=F_{s}(\{R_{h}^{n}\})+F_{s}(\{t_{h}^{n}\})+\sum_{n=1}^{N}E_{mask}^{n}$ (6)
$E_{mask}^{n}$ $=|\mathcal{C}_{h}^{n}\otimes(\mathcal{M}_{h}^{n}-\Pi(V_{h:c}^{n}))|^{2}_{2}$ (7)
where $\Pi(\cdot)$ is the differentiable projection function [24, 34] and $\mathcal{M}_{h}^{n}$ is the hand mask. $\mathcal{C}_{h}^{n}$ is the occlusion-aware mask as in [19, 54] which only computes the error within regions of the hand that are not occluded by the grasped object $\mathcal{C}_{h}^{n}$, set to 1 for the hand and background, and 0 for the object.
4.4 Optimising Object Relative Pose
After refining the hand wrist pose $T_{h2c}^{n}$ and finger articulation $\theta^{n}$ in Sec 4.3, we keep these fixed, and optimise the object relative pose parameters ($P_{o2h}$, $P_{o2\phi}$, $\{R_{\phi}^{n}\}$) and scale $s$ by minimising the following objective function:
$E(P_{o2h},P_{o2\phi},\{R_{\phi}^{n}\},s;\{\theta^{n}\},\{T_{h2c}^{n}\})=\sum\limits_{n=1}^{N}\lambda_{1}E_{mask}^{n}+\lambda_{2}E_{ins}^{n}+\lambda_{3}E_{clo}^{n}$
The term $E_{mask}$ focuses on estimating a reconstruction that best matches the 2D projections of the object masks throughout the sequence. We measure the error via sum of pixel differences:
$E_{mask}^{n}=|\mathcal{C}_{o}^{n}\otimes(\mathcal{M}_{o}^{n}-{\Pi(V_{o:c}^{n})})|^{2}_{2}$ (9)
where $\mathcal{M}_{o}^{n}$ is the object mask and $\Pi(\cdot)$ is the differentiable projection function. Similar to our optimisation for hand vertices using the masks, we wish not to penalise parts of the object occluded by the hand. Accordingly, we define the mask $\mathcal{C}_{o}^{n}$ such that $\mathcal{C}_{o}^{n}$ is 1 for the object and the background, and 0 for the hand.
Characteristics Stats Labels
Dataset Year In-the-wild Funct. Intent Ego #Env #Sub #Cat #Inst #Seq
Pose GT
Stable Grasp
FPHA [15] 2018
3 6 4 4 1,175
3D
HO3D [17] 2020
1 10 10 10 65
3D
✓(part)
ContactPose [2] 2020
1 50 25 25 2,306
3D
GRAB [41] 2020
1 10 51 51 1,334
3D
H2O [26] 2021
3 4 8 8 24
3D
DexYCB [6] 2021
1 10 20 20 1,000
3D
HOI4D [30] 2022
610 9 20 800 5,000
3D
Assembly101 [38] 2022
1 53 15 15 4,321
3D $*{}^{*}$
OakInk [49] 2022
1 12 32 100 1,356
3D
ARCTIC [13] 2023
1 ${}^{\dagger}$ 11 11 339
3D
ARCTIC-Grasps (ours) 2023
1 9 11 11 1,303
3D
Core50 [31] 2017
11 - 10 50 550
2D Mask
MOW [5, 33] 2021
500 500 121 500 500
EPIC-Grasps (ours) 2023
141 31 9 $\sim$390 2,431
2D Mask
Table 1: Dataset Comparison: $*{}^{*}$: cannot be used for hand-object reconstruction as object poses or segments are not provided. ${}^{\dagger}$: subjects in the releasted train/val set. Two subjects are in the held out test set.
Similar to [50, 5, 19], we use the physical heuristics $E_{ins}$, which pushes the object out of the penetrating region against the hand and a balancing loss $E_{clo}$ which pulls the object to touch these contact regions. Please refer to the appendix for more details of the physical heuristics.
The latent axis $P_{o2\phi}$ is initialised at the origin in the object coordinate system and initially points in the $+z$ direction; the $R_{\phi}^{n}$ are initialised to 0. The initialisation of $P_{o2h}$ is described in the implementation details.
4.5 Hand-Object Pose Refinement
In Sec 4.3 the initial hand mesh was refined without the knowledge of the in-hand object. After the optimisation of $T_{o2h}^{n}$, we keep the object relative poses fixed and perform two steps of refinement in an object-aware manner.
Refining hand-object as a single mesh. We glue the per-frame pair of meshes of the hand and the object into a single mesh, i.e. $V_{ho:h}$ as a concatenation of $V_{h:h}$ and $V_{o:h}$. This step refines the hand pose $T_{h2c}^{n}=[R_{h}^{n}|t_{h}^{n}]$ so that the combined hand-object mesh aligns better with the combined hand-object mask (Fig 3 right-most column). We then optimise the sum of pixel differences and the smoothness of the object in the camera coordinate system $V_{o:c}^{n}$:
$E(\{R_{h}^{n}\},\{t_{h}^{n}\})$ $=F_{s}(\{V_{o:c}^{n}\})+\sum_{n=1}^{N}E_{mask}^{n}$ (10)
$E_{mask}^{n}$ $=|\mathcal{C}_{ho}^{n}\otimes(\mathcal{M}_{ho}^{n}-{\Pi(V_{ho:c}^{n})})|^{2}_{2}$ (11)
where $\mathcal{C}_{ho}^{n}$ is 1 for the hand-object and the background, and 0 for other objects; $M_{ho}^{n}$ is the combined hand-object mask.
Refining finger articulation. As the object has been introduced to the hand, we refine finger articulations such that the hand and object are in plausible contact, using physical constraints $E_{ins}$ and $E_{clo}$ introduced in Sec 4.4:
$E(\{\theta^{n}\})=F_{s}(\{\theta^{n}\})+\sum_{n=1}^{N}E_{ins}^{n}+E_{clo}^{n}$ (12)
As summarised in Fig 3, we perform three stages of optimisation. The first stage refines FrankMocap output using ground-truth hand masks. The second optimises the in-hand object pose using known hand poses and ground-truth object masks. The third refines the hand pose and the finger articulation in an object-aware manner.
5 EPIC-Grasps Annotations
With the definition of stable grasp finalised on in-lab videos of the ARCTIC dataset, we now use this to annotate the temporal extents of stable grasps in unscripted egocentric recordings capturing in-the-wild hand-object stable grasps. This offers a dataset distinct from prior works, where frame-level labels [49, 20, 5], 3D supervision [2, 41, 53] or recordings specifically collected to evaluate grasps with no underlying action [17, 26, 6] are typically used. Instead, we aim to find stable grasps from within egocentric videos recorded of daily actions in an unscripted manner.
We next detail our pipeline to annotate EPIC-Grasps:
1. Identifying candidate clips to annotate. The ultimate goal of hand-object reconstruction is to operate for any rigid or dynamic objects, including novel previously unknown objects. However, current approaches for reconstruction of unknown objects are still in their infancy [51, 52, 39, 22, 40]. We thus restrict our first-attempt at reconstruction in-the-wild to objects for which a category CAD model can be acquired. Note that this is distinct from instance CAD models in ARCTIC-Grasps, for example, the general CAD model of a bottle might not exactly match all bottles in daily life.
We exclude very small objects and shortlist 9 object categories frequently used in kitchens: plate, bowl, bottle, cup, mug, can, pan, saucepan, glass [^1] For the object mesh, we made the CAD model in Blender [8] as the common shape of each category..
We begin with the frame-level hand-object interaction annotations from VISOR [12], in which we identified clips where the a hand is in contact with one of the categories above. The annotations also inform us which hand is in contact with the object (i.e. left or right). We then use the fine-grained action labels available from EPIC-KITCHENS [11, 10] to locate the initial temporal extent for actions involving any of these categories. This resulted in 5516 in-action candidate segments.
2. Annotating Stable Grasps. Two annotators were asked to label the start-and-end frames following the stable grasp definition in Sec 3. We discard segments when (i) a segment does not contain any stable grasp (ii) both the hand and object are out-of-view during the sequence (iii) the object does not match the category CAD model specified.
3. 2D Hand and Object Segmentation Annotations. In total, we labelled 2431 video clips of stable gasps from 141 distinct videos in 31 kitchens [11]. For each clip, we have a start and end time of the stable grasp, as well as 319,661 segmentation masks for the hand and the object during the stable grasp from the dense VISOR annotations [12]. Of these, 1446 contain left hand stable grasps and 985 contain right hand stable grasps. Note that the majority of subjects recording the EPIC-KITCHENS dataset are right-handed—which implies the left hand is more frequently used to steady the object during manipulation. Compared to previous efforts [17, 30, 26, 6, 40], most clips in EPIC-Grasps contains both hands in view manipulating the same or different objects.
Table 1 compares our dataset with currently available and regularly used datasets for hand-object reconstruction. In the first section, we show datasets collected with 3D Ground-Truth. These datasets are limited in the environmental setup to allow for multi-camera views or motion capture tracking. This limits the appearance diversity and interactions are not functional. In Table 1 we also report the functional intent of various dataset (i.e. object is used with intent). our dataset is the first to capture in-the-wild egocentric videos with functional intent. Of particular note is the HOI4D [30] where diversity is targeted with a large number of indoor rooms. However, this is a constructed setup, where objects are selected from a given pool. In EPIC-Grasps, subjects are recording in their actual (personal) kitchens and objects.
6 Experiments
The experiment section is structured as follows: in Sec 6.1 and Sec 6.2 we describe the implementation details of our method and the quantitative metrics respectively. Sec 6.3 and Sec 6.4 report results of our method and baselines on ARCTIC-Grasps and EPIC-Grasps, respectively. Finally Sec 6.5 conducts ablation studies on our method.
6.1 Implementation Details
Finding stable grasps on ARCTIC. We measure the object in-contact area using set of vertices. As individual vertices are sensitive to noise, we partition vertices into fragments using the farthest-point-sampling method, and measure the area using unions of fragments. We set the number of fragments to 0.01 of the number of vertices. As our pipeline assumes rigid object, we exclude segments where the hand is changing the object’s articulation.
Optimisation Pipeline. We follow the convention in HOMan [19] which sparsely and linearly samples 30 frames from the sequence for optimisation and evaluation. We use 50 initialisations of object rotation and 1 global translation for each object [^2] We use prior knowledge of grasps from ground-truth of ARCTIC to find these initialisations. The optimisation process takes 1.5 minutes on a 2080Ti for one 30-frame sequence on average.
[…]
Please refer to appendix for further implementation details.
6.2 Baselines and Quantitative Metrics
We use 4 baselines to compare to:
Single Frame HOMan [19] Static Dynamic 1-DoF
IOU SCA-IOU ADD SCA-ADD IOU SCA-IOU ADD SCA-ADD IOU SCA-IOU ADD SCA-ADD IOU SCA-IOU ADD SCA-ADD IOU SCA-IOU ADD SCA-ADD
box 93.2 17.5 9.4 3.1 83.5 30.9 33.3 16.9 90.9 67.3 37.7 28.7 95.2 41.7 57.2 25.4 94.2 56.7 39.1 23.6
capsulemachine 79.3 17.8 8.4 3.3 84.0 34.4 42.1 24.5 79.8 34.9 48.4 34.1 88.8 41.6 53.7 29.2 86.1 51.4 62.1 41.2
espressomachine 82.8 19.1 22.8 8.3 84.2 36.9 44.6 28.8 82.6 49.2 48.5 36.2 89.6 43.1 66.3 34.5 87.4 54.9 54.5 35.6
ketchup 81.3 29.3 30.2 16.4 72.0 18.0 15.1 9.3 75.3 32.7 51.9 37.8 89.4 56.0 62.3 40.2 84.0 55.4 64.2 45.0
laptop 87.8 24.9 16.7 7.0 82.1 35.3 43.8 28.0 86.0 62.0 45.1 37.1 93.3 50.2 63.2 34.6 91.4 68.0 54.9 40.8
microwave 85.7 17.1 29.5 8.0 87.0 36.1 56.2 27.3 85.4 51.5 50.9 35.3 90.3 35.2 53.6 23.0 89.0 52.5 55.4 34.8
mixer 79.4 12.1 14.8 4.9 85.3 34.3 45.1 26.7 79.3 37.9 48.4 32.2 89.1 38.4 59.0 29.0 87.0 52.0 62.3 39.6
notebook 87.2 25.0 6.6 2.8 85.7 38.7 33.8 20.8 84.5 57.8 43.0 33.2 93.5 52.8 63.6 35.0 90.9 65.5 55.6 38.7
phone 82.9 25.7 15.1 6.8 80.9 39.5 28.1 19.9 77.8 36.7 39.7 29.2 91.1 58.8 37.7 23.5 88.1 61.9 54.1 37.6
scissors 48.5 0.0 14.0 9.3 40.9 5.2 7.0 5.2 48.2 0.0 47.4 36.3 73.9 18.2 56.1 38.9 62.4 2.7 68.4 51.8
waffleiron 85.7 13.1 3.1 0.9 86.5 36.3 45.0 26.2 84.7 48.3 59.5 39.2 92.7 44.3 81.7 38.1 91.2 55.9 69.5 43.7
All 83.3 19.5 15.0 6.0 81.4 33.1 37.1 22.0 81.4 46.6 46.9 34.2 90.8 45.5 59.6 31.5 88.1 55.7 57.3 38.5
Table 2: Results on ARCTIC-Grasps. Green shows the best performing method per metric and yellow shows second best.
  • HOMan [19]—a common baseline in other works [33, 52] that progressively optimises the object.
  • Single Frame Single-frame optimisation without utilising the knowledge of stable grasps.
  • Static A version of our method where objects are not allowed any motion when within the stable grasp, minimising the overall rotation and translation of the object.
  • Dynamic A variation of our method where objects are allowed to move freely within 6-DoF.
We propose quantitative metrics to measure the correctness of the predicted object poses as well as correctness of the stable grasp. We utilise our stable grasp annotations so as to correctly assess reconstruction with/without 3D ground truth. Specifically, given the knowledge that stable grasps maintain a consistent contact area between the hand and the object throughout the sequence, we measure this contact area over time. However, a reconstruction can maintain a stable contact area but be completely erroneous. We thus propose to combine a measure of correctness of reconstruction with the stability of the contact area to produce a robust metric. We explain this next.
Average Distance (ADD %). Following [25, 47, 21], we measure the distance of corresponding vertices between GT and predicted object vertices in the hand coordinate, and average this distance over vertices and frames. ADD is 1 for a sequence if the average distance is less than 10% of the object’s diameter, and 0 otherwise.
Average Stable Contact Area at ADD Success (SCA-ADD %). When a pose is considered correct for a sequence, i.e. ADD is 1, we measure the Stable Contact Area across the sequence, defined as the average IOU of surface in contact area between each pair of frames 3. SCA-ADD is set to 0 when ADD is 0. We average over all examples.
Intersection-over-Union (IOU %). We use IOU as a proxy of pose accuracy when 3D GT is not available. We measure the Intersection-over-Union between the ground-truth mask and the rendered mask for the object in camera view. We report average IOU across all frames. Due to occlusion with other components in the scene, only the non-occluded area of the rendered projection is used.
Average Stable Contact Area at high IOU (SCA-IOU %). Using IOU as proxy for correct reconstructions, we analogously report SCA when IOU is more than certain thresholds. We use 80% as threshold on ARCTIC-Grasps and report both 80% and 60% on EPIC-Grasps. SCA-IOU is set to 0 when IOU is below the threshold. We report average SCA-IOU.
The first 2 metrics, ADD and SCA-ADD, are evaluated when 3D ground truths are available—e.g. ARCTIC-Grasps dataset; and the other two are the corresponding 2D proxy, which we use for EPIC-Grasps.
6.3 Results on ARCTIC-Grasps
Table 2 compares results on ARCTIC-Grasps dataset using all the metrics. We focus solely on the object stable grasp optimisation in the ARCTIC-Grasps dataset, thus using ground-truth hand mesh and camera parameters. We also use a known scale for the object. We thus do not use the hand refinement stages to report these results.
While Dynamic and 1-DoF both have significantly higher ADD results, 1-DoF maintains the best SCA when grasps are successfully reconstructed (high SCA-ADD). While Static has considerably high SCA, its IOU and ADD are significantly lower than 1-DoF in every category. HOMan only performs best on the “microwave” object sequences. Single Frame performs worst in every case.
6.4 Results on EPIC-Grasps
Table 3 compares results on EPIC-Grasps dataset using proxy metrics IOU and SCA-IOU. We also report SCA-IOU at IOU threshold 0.6 since the overall IOU for are lower. We evaluate two best variations based on 3D performance on ARCTIC-Grasps: Dynamic and 1-DoF. 1-DoF achieves the best SCA-IOU metric for both thresholds for every object category and overall.
HOMan [19] Dynamic 1-DoF
IOU SCA@0.8 SCA@0.6 IOU SCA@0.8 SCA@0.6 IOU SCA@0.8 SCA@0.6
bottle 56.7 3.4 5.8 64.8 7.0 20.4 62.7 10.2 33.1
bowl 54.4 1.1 1.6 58.5 7.6 12.4 57.2 12.2 22.2
can 48.3 3.4 5.1 52.4 10.1 15.4 52.4 12.7 24.6
cup 56.9 5.5 7.8 61.4 6.0 24.2 60.3 10.1 36.8
glass 55.4 3.0 4.2 60.2 8.2 21.7 57.5 10.2 29.7
mug 59.4 3.9 6.6 59.0 5.0 18.0 60.1 6.7 28.7
pan 48.4 0.5 1.9 47.0 2.5 6.9 47.5 4.4 13.5
plate 61.1 1.0 1.6 67.1 12.9 19.3 66.6 18.7 30.3
saucepan 51.1 0.4 2.4 57.2 3.4 11.5 55.9 9.0 19.7
All 54.9 1.9 3.3 58.6 7.2 15.5 57.8 10.8 25.2
Table 3: Results on EPIC-Grasps. Green for best method and yellow shows second best.
Figure 4 shows reconstruction results of our 1-DoF approach on the EPIC-Grasps dataset. We showcase three examples per category. From the images, the challenge in EPIC-Grasps dataset is evident with the hand and object occupying a small proportion of the image. In most examples, the palm and fingers are concealed. Importantly, all grasps are functional – e.g. picking a cup from the drainer.
Figure 4: Qualitative Results on EPIC-Grasps (3 examples per category); input frames (top), projected reconstruction results (bottom) and reconstruction in rotated views (right).
Figure 5 shows comparative results. Despite acceptable overlay, HOMan does not maintain the stable grasp and the object is in fact sliding in the hand over time. The Dynamic baseline has significantly lower SCA with the contact area drifting over time. Our proposed 1-DoF method is able to reconstruct the stable grasp as it benefits from the constrained relative pose assumption.
Figure 5: Qualitative comparison between 1-DoF, Dynamic and HOMan [19]. We show input frames (top), grasp reconstruction results and the corresponding SCA.
While results in-the-wild are very promising, our pipeline relies on FrankMocap as a first stage. Despite the robustness incorporated by the multiple-view joint optimisation, our method fails when the majority of hand poses are incorrect (see Fig 6).
6.5 Ablation Studies
What happens outside the stable grasp boundary? In Table 4 we show results when we remove the knowledge of the grasp start-end time. We run this ablation on a random subset of ARCTIC-Grasps dataset of 100 sequences, for tractable processing. We extend the start-end time of the boundary by 0.5s on both sides.
Clearly, all methods show a drop in performance, as some of these frames might not even have the hand in grasp of the object. Dynamic has the least drop as it does not have a constraint to optimise.
Static Dynamic 1-DoF
ADD SCA-ADD ADD SCA-ADD ADD SCA-ADD
Within 45.0 31.4 55.0 26.7 54.0 33.3
Outside 11.0 6.9 46.0 15.1 27.0 12.9
Table 4: Within/Outside grasps on ARCTIC-Grasps Sequences.
Ablation on Refinement Table 5 presents the impact of performing two refinements on our method. The final refinement stage alone leads to a small improvement; when using both refinement stages, this leads to an average increase of 3.6% in IOU and 7.4% in SCA@0.6.
Init Refine HO Refine IOU SCA@0.8 SCA@0.6
54.1 6.4 18.9
54.7 9.2 23.4
57.8 10.8 25.2
Table 5: Impact of refinement stages on EPIC-Grasps.
Figure 6: Failure Case in EPIC-GRASPS due to FrankMocap.
7 Conclusion
In this paper we build a challenging in-the-wild egocentric videos dataset for hand-object reconstruction. The dataset contains 2D masks as pseudo ground-truths and manual temporal labels of stable grasps. By focusing on stable grasps, we can jointly optimise multiple frames given the knowledge of how objects move relative to the hand within grasp. We evaluate the method’s ability to achieve a stable grasp without 3D ground truth by approximating the 2D IOU of object masks along with measuring the stable contact area over time.
By investigation stable grasps, we hope to encourage more methods to quantitatively evaluate methods in-the-wild. An obvious limitation of our approach is its reliance on the knowledge of the category’s CAD model. Current CAD-agnostic methods struggle in-the-wild – we showcase failures in appendix [51]. However, removing the limitation of CAD model knowledge is a clear future direction.
Acknowledgements. Research is supported by EPSRC UMPIRE (EP/T004991/1). Z. Zhu is supported by UoB-CSC Scholarship.
Appendix
This appendix is structured as follows: in Sec. A we presents statistics of EPIC-Grasps annotation. Sec. B showcases failure of current CAD-agnostic approach on EPIC-Grasps dataset. Sec. C provides further implementation details.
Appendix A Statistics of EPIC-Grasps annotation
In Table 6, we present average, standard deviation and maximum clip length for each of our categories as well as total number of video clips per category. Note that grasping some categories (e.g. Pan) can be significantly longer in duration than others (e.g. Mug).
length (sec)
category avg std. max #-clips
Bottle 4.25 4.70 29.40 311
Bowl 3.59 5.53 43.36 421
Can 3.50 4.11 22.76 105
Cup 2.79 3.20 19.72 125
Glass 2.45 3.73 34.72 226
Mug 2.40 3.34 29.74 167
Pan 6.39 9.36 91.28 526
Plate 2.77 4.49 54.60 473
Saucepan 5.73 10.08 76.10 77
All 3.95 6.30 91.28 2431
Table 6: EPIC-Grasps sequences statistics
Appendix B Evaluating CAD-agnostic method Ihoi
In our method, we assume knowledge of the CAD model. We also explored the latest works that attempt hand grasps without such knowledge. In this section, we showcase these models to be unusable for our in-the-wild scenes through three examplese, qualitatively.
Figure 7: In-the-wild test of Ihoi [51]
The recent shape reconstruction method proposed in [22] requires careful scan of the underlying in-hand object hence is not suitable for testing in our case. Therefore, we evaluate the state-of-the-art image-based CAD-agnostic method Ihoi [51] on the EPIC-Grasps dataset, using a model trained on HO3D [17] images that contain the similar categories: mug and bottle – i.e. the method is aware of these shapes during training. Figure 7 shows Ihoi is unable to generate plausible object meshes.
Appendix C Further Implementation Details
Physical Heuristics In Eq (8) of the main paper we note our usage of physical attraction and repulsion losses $E_{ins}$ and $E_{clo}$.
The term $E_{ins}$ ensures all object vertices are located inside the contact surface of the hand (Fig 8). For each $v_{o}\in V_{o}$, we locate the nearest vertex in hand contact regions, and compute distance along the surface normal of this hand vertex. Object vertices that penetrate into the contact surface will have negative values. We maximise those negative values, truncating the positive ones:
$E_{ins}^{n}$ $=\sum_{v_{o}\in V_{o}}-1*\min(d_{v},0)$ (13)
$d_{v}$ $=\langle v_{o}-v^{*}_{h},n^{*}_{h}\rangle$ (14)
where $v^{*}_{h}$ is the corresponding nearest vertex on the hand and $n^{*}_{h}$ is the surface normal of $v^{*}_{h}$.
In addition to $E_{ins}$, which pushes the object out of the penetrating region against the hand, we use a balancing loss $E_{clo}$ which pulls the object to touch these contact regions. We here focus on the eight contact regions showcased in Fig 8. For each contact region with hand vertices $\{v_{h}\}_{C}$, the region-to-object distance is defined as the minimum distance of all $(v_{h},v_{o})$ pairs. We use 5 finger tip regions and minimise the average of these region-to-object distances.
$E_{clo}^{n}$ $=\frac{1}{5}\sum_{C}d(\{v_{h}\}_{C},V_{o})$ (15)
$d(\{v_{h}\}_{C},V_{o})$ $=\min_{v_{h}\in\{v_{h}\}_{C},v_{o}\in V_{o}}\langle v_{h}-v_{o},n_{o}\rangle$ (16)
where $n_{o}$ is the surface normal of $v_{o}$.
Figure 8: Eight contact regions: five finger tips + three palm areas. The contact regions serve two purposes: bounding the object inside and attracting the object closer to these regions.
Initialisation of poses on EPIC-Grasps For EPIC-Grasps dataset, we manually set initial object relative poses to the common poses of each category, results in 80 total initial poses for 9 object categories.